Calculate the wavelength of the photon emitted when an electron makes a transition from n=6 to n=3. You can make use of the following constants: h=6.626×10^−34 J⋅s c=2.998×10^8 m/s 1 m=10^9 nm

Express your answer to four significant figures and include the appropriate units.
Answer #1

Given that

initial state, ni = 6

Final state, nf = 3

From Rydberg equation

1/ = RH ( 1/nf2 - 1/ni2)

Where = wavelength

and RH = Rydberg constant = 1.097 *107 m-1

1/ = (1.097 *107 m-1) ( 1/(3)2 - 1/(6)2 ) = (1.097 *107 m-1)(1/9 - 1/36) = 9.14 *105 m-1

= 1094 nm